Find the length of a ramp that is 8 feet long and 6 feet high
Answer by smmoore(10)

this is a simple problem that requires you to use the pythagorean theorem.
to find the length of the ramp you need to set up a right triangle with the height being 8 and the length/base as 6. then the hypotenuse is the missing length.
the pythagorean theorem is a^2+b^2=c^2
then you plug in 8 and 6 in for a and b, because c is always the hypotenuse. now it really doesn't matter which leg of the right triangle is labeled a and b
I will put 8 in for a and 6 in for b.
8^2+6^2=c^2 is what we now have
64+36=c^2
100=c^2 .... is what we now have
then we must now square root both sides of what we have now
therefore we now have 10=c, and since c is the missing length in feet, as stated in the problem, the ramp is 10 feet in length.
ramp= 10 feet in length.